OpenStudy (anonymous):

used the illustration below to calculate the circumfrence of the cirlcle. A. 17.5 cm B. 15.7 cm C. 31.4 cm D. 78.5 cm

OpenStudy (mertsj):

\[C=2\pi r\]

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry ignore that line on the right

OpenStudy (johnweldon1993):

well it looks as though the hypotenuse of that right triangle....would be the diameter of the circle right?...so how do we find that?

OpenStudy (anonymous):

a2+b2=c2

OpenStudy (johnweldon1993):

yes! by using the pythagorean theorem of course! A² + B² = C² 3² + 4² = C² C = √(3² + 4²) so C =?

OpenStudy (anonymous):

a2+b2+c2 right?

OpenStudy (anonymous):

is that right?

OpenStudy (johnweldon1993):

A² + B² "=" C² this is how you solve for a missing side of a right triangle.....for example the hypotenuse...is equal to the square root of the other 2 sides squared so C = √A² + B²

OpenStudy (johnweldon1993):

now you already have the 2 other sides....you just want the hypotenuse so plug in your numbers C = √(3² + 4²) so C = ?

OpenStudy (anonymous):

c=5

OpenStudy (johnweldon1993):

right! so now you have...

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

as you can see...that side (5cm) is acting as the diameter of the circle... now if we have the diameter.....how do we find the radius??

OpenStudy (anonymous):

that line way to the right was an accident when drawing it lol

OpenStudy (anonymous):

anyways I got my answe as 15.7

OpenStudy (johnweldon1993):

lol yeah I guessed that lol.....and yes very good :))

OpenStudy (anonymous):

Thanks!

OpenStudy (johnweldon1993):

anytime! :)

OpenStudy (johnweldon1993):

really!???

OpenStudy (johnweldon1993):

lol please correct me then! I really should brush up on my geometry!

OpenStudy (johnweldon1993):

are you sure? because trust me geometry wasn't my best subject lol....so there is a definite possibility....but I felt pretty good about this one :)